# Modified gravity interpretation of the evolving dark energy in light of DESI data

### v1

@@ IV Conclusions
Taking this preference seriously, we study a physical model which can explain DESI’s preference for evolving dark energy in both background and perturbation dynamics. We explore a general Horndeski scalar-tensor gravity theory, which allows the dark energy equation of state to evolve across the phantom divide, $w=-1$. Utilizing the EFT-based framework, which offers a largely model-independent description of modified gravity effects, we derive constraints for $w_{0}$ and $w_{a}$ parameters, showing good consistency with the $w_{0}w_{a}{\rm CDM}$ model. The modified gravity model indicates a preference over $\Lambda{\rm CDM}$ at the $2.4\sigma$ level. We conclude that modified gravity offers a viable physical explanation for DESI’s preference for evolving dark energy.
@@ Acknowledgments
We are grateful to Emilio Bellini for fruitful discussions.
@@ Appendix A Full Parameter Constraints
Model/Dataset $H_{0}$ $\Omega_{\mathrm{m}}$ $\sigma_{8}$ $w_{0}$ $w_{a}$ $\alpha_{B}$ $\alpha_{M}$
@@
CMB $69.22_{-5.10}^{+6.39}$ $0.302_{-0.063}^{+0.032}$ $0.850_{-0.059}^{+0.062}$ $-0.931_{-0.478}^{+0.315}$ $-0.43_{-0.64}^{+1.16}$ $0.84_{-0.49}^{+0.45}$ $0.87_{-1.50}^{+0.73}$
CMB+DESI $66.67_{-3.30}^{+2.98}$ $0.321_{-0.034}^{+0.028}$ $0.824_{-0.043}^{+0.028}$ $-0.708_{-0.338}^{+0.270}$ $-0.97_{-0.62}^{+0.97}$ $0.84_{-0.41}^{+0.32}$ $0.62_{-1.29}^{+0.54}$
CMB+DESI+SN $67.81_{-0.72}^{+0.73}$ $0.308_{-0.007}^{+0.007}$ $0.837_{-0.021}^{+0.015}$ $-0.856_{-0.064}^{+0.060}$ $-0.53_{-0.26}^{+0.28}$ $0.82_{-0.46}^{+0.31}$ $0.88_{-1.28}^{+0.75}$
Fig. 5 illustrates the degeneracies between various parameters present in the CMB data. The DESI distance measurements significantly improve constraints on the $\Lambda{\rm CDM}$ parameters, although the degeneracies among the $w_{0}$, $w_{a}$, and other parameters persist. The addition of SN Ia data completely breaks the degeneracy in the $w_{0}-w_{a}$ plane, thereby refining the estimates of other parameters.

### v2

@@ IV Conclusions
Taking this preference seriously, we study a physical model which can explain DESI’s preference for evolving dark energy in both background and perturbation dynamics. We explore a general Horndeski scalar-tensor gravity theory, which allows the dark energy equation of state to evolve across the phantom divide, $w=-1$. Utilizing the EFT-based framework, which offers a largely model-independent description of modified gravity effects, we derive constraints for $w_{0}$ and $w_{a}$ parameters, showing good consistency with the $w_{0}w_{a}{\rm CDM}$ model. The overall results show a $2.4\sigma$ discrepancy with $\Lambda{\rm CDM}$, as estimated from the best-fit $\chi^{2}$ values. We conclude that modified gravity offers a viable physical explanation for DESI’s preference for evolving dark energy.
@@ Acknowledgments
We are grateful to Emilio Bellini for fruitful discussions. Numerical calculations have been performed with the Baobab high-performance computing cluster at the University of Geneva.
@@ Appendix A Full Parameter Constraints
Model/Dataset $H_{0}$ $\Omega_{\mathrm{m}}$ $\sigma_{8}$ $w_{0}$ $w_{a}$ $c_{B}$ $c_{M}$
@@
CMB $68.15_{-6.75}^{+6.63}$ $0.314_{-0.075}^{+0.046}$ $0.854_{-0.070}^{+0.068}$ $-0.857_{-0.447}^{+0.323}$ $-0.42_{-0.61}^{+1.22}$ $0.92_{-0.45}^{+0.42}$ $0.67_{-1.33}^{+0.65}$
CMB+DESI $66.70_{-3.38}^{+3.16}$ $0.321_{-0.035}^{+0.028}$ $0.825_{-0.044}^{+0.032}$ $-0.721_{-0.355}^{+0.267}$ $-0.91_{-0.60}^{+1.0}$ $0.85_{-0.40}^{+0.34}$ $0.69_{-1.34}^{+0.53}$
CMB+DESI+SN $67.78_{-0.71}^{+0.73}$ $0.309_{-0.007}^{+0.007}$ $0.837_{-0.022}^{+0.015}$ $-0.859_{-0.062}^{+0.062}$ $-0.52_{-0.25}^{+0.28}$ $0.84_{-0.46}^{+0.31}$ $0.89_{-1.31}^{+0.78}$
Fig. 5 illustrates the degeneracies between various parameters present in the CMB data. The DESI distance measurements significantly improve constraints on the $\Lambda{\rm CDM}$ parameters, although the degeneracies among the $w_{0}$, $w_{a}$, and other parameters persist. The addition of SN Ia data breaks the degeneracy in the $w_{0}-w_{a}$ plane, thereby refining the estimates of cosmological parameters.